Circular No. 8459

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2004gw AND 2004gx Further to IAUC 8435, T. Puckett and R. Ireland report the discovery

of an apparent supernova (mag 17.9) on an unfiltered CCD frame taken with the 0.60-m automated supernova patrol telescope on 2004 Dec. 27.14 UT (and confirmed on frames taken on Dec. 29.11 at mag 17.7). The new object is located at α = 5^{h}08^{m}48^{s}.41, δ = +62^{o}26′20′′.7 (equinox 2000.0), which is 17′′.7 east and 14′′.2 south of the center of PGC 16812. The object is not present on images taken by Puckett on Sept. 10 and Nov. 9 (limiting mag ∼ 20.0).

D. Rich, Hampden, ME, reports his discovery of an apparent supernova (mag ∼ 17.6) on unfiltered CCD frames taken with a 0.31-m reflector on 2004 Dec. 30.96 and 2005 Jan. 1.93 UT. SN 2004gx is located at α = 23^{h}33^{m}30^{s}.80, δ = +24^{o}01′07′′.0 (equinox 2000.0), which is ∼ 7′′.0 west and 5′′.4 south of the center of UGC 12663. Nothing is visible at this location on Palomar Sky Survey plates (limiting red mag 20.1, limiting blue mag 19.7) or on CCD frame taken by Rich on 2004 Oct. 21.01 and Nov. 20.07 (limiting mag ∼ 18.6).

SUPERNOVA 2005A IN NGC 958 Further to IAUC 8453, J. Graham and W. Li report the LOSS discovery

of an apparent supernova (mag ∼ 17.1) on unfiltered KAIT images taken on Jan. 5.26 and 6.25 UT. SN 2005A is located at α = 2^{h}30^{m}43^{s}.25, δ = −2^{o}56′19′′.8 (equinox 2000.0), which is 8′′.2 east and 0′′.7 north of the nucleus of NGC 958. A KAIT image taken on 2004 Dec. 17.25 showed nothing at this position (limiting mag ∼ 19.0).

SGR 1806–20 P. Chandra, Tata Institute of Fundamental Research, Mumbai, reports

the low-frequency radio detection of the soft γ-ray repeater SGR 1806–20 (cf. IAUC 5875) with the Giant Metrewave Radio Telescope on Jan. 4.409 UT. The flux density at 608 MHz was 198.9 ± 11.9 mJy, while at 240 MHz it was 242.6 ± 14.9 mJy. The power-law index between the two frequencies is −0.21 ± 0.09, which is substantially flatter than that reported for the higher-frequency VLA observation on Jan. 4.59 by Cameron and Kulkarni (GCN 2934; α = −0.74 at 1.4 GHz) or −0.61 ± 0.06 reported by Gaensler et al. at the same time and frequency (GCN 2933).

2005 January 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green